What is the pH of the solution given that the concentration of H2PO4- is 0.1 and that of HPO42- is 0.066?

1
Expert's answer
2021-09-21T01:55:25-0400

Solution:

The Henderson–Hasselbalch equation can be used.

The formula for the Henderson–Hasselbalch equation is: 

pH = pKa + log([A] / [HA]),

where [A] and [HA] are concentrations of the conjugate base and starting acid.


H2PO4 – starting acid

HPO42–– conjugate base

Thus,

[A] = 0.066 M

[HA] = 0.100 M


H2PO4 ⇔ HPO42– + H+, pKa = 7.2


Hence,

pH = 7.2 + log(0.066 / 0.100) = 7.2 + (-0.18) = 7.02

pH = 7.02


Answer: pH = 7.02